# Target-Date Funds: A State-of-the-Art Review with Policy Applications to Chile’s Pension Reform

### v1

@@ 3.3 The 2025 Chilean Pension Reform
The total pension contribution rate will gradually increase from 11.5% to 18.5% of taxable income, entirely paid by employers over a nine-year period. The additional contributions will fund individual pension accounts (4.5%), gender equity initiatives (1%), and a new fund for social protection (1.5%). The reform also introduces the Seguro Social Previsional and the Fondo Autónomo de Protección Previsional (FAPP), which aim to address gender disparities in pensions and manage long-term resources for retirees.
To support vulnerable groups, the Pensión Garantizada Universal (PGU) will gradually increase to CLP \$250,000, with expanded coverage for additional vulnerable populations.
@@
In addition to the structural changes, the reform introduces a new system of variable fees and penalties tied to fund performance. Under this system, if the performance of a Generational Fund exceeds its benchmark plus the maximum allowable deviation, calculated over a 36-month period, the AFP will be entitled to an additional fee, capped at 15% of the average commissions charged by all AFPs. Conversely, if the fund’s performance falls below the benchmark minus that allowable deviation, the AFP will be required to compensate the fund, with the same 15% cap applied to the amount contributed. This creates a performance band around the benchmark within which no reward or penalty applies. The performance-based fee structure is designed to align the interests of AFPs with the long-term success of participants’ investments, incentivizing stronger fund management while holding AFPs accountable for underperformance.
4 Considerations in Designing the Glide Path and Investment Regime

### v2

@@ 3.3 The 2025 Chilean Pension Reform
According to the pension reform, a fundamental point to emphasize is the long-term increase in the total mandatory savings contribution rate directed to individual accounts, rising from the current 10% to an aggregate 16% of income. This significant increase occurs as the existing 10% contribution will be supplemented by an additional 6% mandatory savings contribution funded by the employer. Although the total new mandatory contribution is set at 18.5%, 2.5% will be allocated to the Social Insurance fund. The remaining 6% consists of a 4.5% direct employer contribution to the individual account and a 1.5% transitional deferred contribution known as the “Contribution with Protected Return” (Aporte con Rentabilidad Protegida), which is also ultimately destined for individual accounts. This 1.5% contribution is transitional for 30 years and will be disbursed upon the worker’s retirement to fund a benefit based on the number of years contributed. Consequently, the most significant long-term implication for individual retirement savings is the elevation of the total mandatory contribution rate to 16%.
To support vulnerable groups, the Pensión Garantizada Universal (PGU), a state-funded, non-contributory pension aimed at ensuring a minimum income for elderly individuals, will gradually increase to CLP \$250,000. In addition to the higher benefit amount, the reform expands the eligibility criteria, allowing broader coverage of vulnerable populations who were previously excluded due to income or other requirements.
@@
In addition to structural changes, the reform modifies the reserve (known as the “encaje”) mechanism that AFPs must maintain. Instead of a fixed percentage of each fund’s assets, the total reserve now corresponds to 30% of the AFP’s commissions collected over the previous 12 months. This total is then proportionally allocated across all funds managed by the AFP, based on their relative size. According to the reform, the purpose of this reserve is to cover any mandatory contributions that the AFP must make to a fund when its monthly return falls below a minimum “performance unit” defined in the Investment Regime. In this way, the reform links fund underperformance to AFP accountability through this reserve mechanism.
4 Considerations for Designing the Glide Path and Investment Regime in the Chilean Context